OpenStudy (anonymous):

Find x,if 2tan^-1 (1/3)+tan^-1 (1/7) = tan^1 x . . ANSWER X=1

OpenStudy (anonymous):

\[2\tan ^{-1} \frac{ 1 }{3 }+\tan ^{-1}\frac{ 1 }{ 7 }=\tan ^{-1}x\]

OpenStudy (anonymous):

I am havin problem in end

OpenStudy (anonymous):

Let\[\theta=\tan ^{-1}x\] \[x=\tan \theta\]

OpenStudy (anonymous):

x=tan[2tan^-1 (1/3) +tan^-1 (1/7)]

OpenStudy (anonymous):

Let A = tan^-1 (1/3) tanA=1/3

OpenStudy (anonymous):

Let B = tan^-1 (1/7) tanB=1/7

OpenStudy (anonymous):

x=tan{tan2A+tanB}

OpenStudy (anonymous):

\[As \tan(a-b) = \frac{ tana+tanb }{ 1-tana*tanb }\]

OpenStudy (harsimran_hs4):

x = tan(2A + B) not x=tan{tan2A+tanB}

OpenStudy (harsimran_hs4):

Glad you figured out your problem :)

OpenStudy (anonymous):

wait I am doing it

OpenStudy (harsimran_hs4):

ok

OpenStudy (anonymous):

what is the formula for tan2A ?

OpenStudy (anonymous):

\[x=\frac{ \tan2A+tanB }{1-\tan2A*tanB }\]

OpenStudy (anonymous):

we know these: tanA=1/3 and tanB=1/7 what is formula for calculating 2A

OpenStudy (harsimran_hs4):

\[\tan (A + B) = \frac{ \tan A + \tan B }{ 1 - (\tan A \tan B) }\] put A = B for tan 2A

OpenStudy (anonymous):

yes I got it "the formula"

OpenStudy (anonymous):

It says:

OpenStudy (anonymous):

\[\tan2 \theta = \frac{ 2\tan \theta }{ 1-\tan^2 \theta }\]

OpenStudy (harsimran_hs4):

yes thats right!!

OpenStudy (anonymous):

tan2A = 2tanA/1-tan^2A

OpenStudy (harsimran_hs4):

yes now you calculate tan2A and tell me ? or if you can then go for answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3/4

OpenStudy (shubhamsrg):

why don;t you simply mug up the formula : tan^-1 a + tan^-1 b = tan^-1 (a+b)/(1-ab)

OpenStudy (harsimran_hs4):

@shubhamsrg let @koli123able concentrate of concepts initially ....my suggestion other what @koli123able feels fine we will tell accordingly

OpenStudy (anonymous):

I got it x=1

OpenStudy (harsimran_hs4):

cool so main mistake was writing x=tan{tan2A+tanB} instead of x = tan(2A + B) so remember this part

OpenStudy (anonymous):

yes thanks

OpenStudy (harsimran_hs4):

:)